# SOLAR 10.7B: Scaling Large Language Models with Simple yet Effective Depth Up-Scaling

### 23 Dec 2023 v1

@@ Appendix A Contributions
  • Innovative LLM Up-Scaling Method: Depth up-scaling is both effective and efficient while also being straightforward to use. The proposed method serves as an easier alternative to other up-scaling methods such as MoE without compromising in efficacy or efficiency,
  • Introduction of the World’s First 10.7 Billion-Parameter Model: SOLAR 10.7B sets a new precedent in the field, demonstrating an unprecedented scale in language model development.
@@
  • Deployment under a Commercially Viable License: The release of SOLAR 10.7B under the Apache 2.0 license allows for commercial use, facilitating the integration of these models into various products and services, thereby bridging the gap between academic research and practical applications.
@@ B.2 Mixture of Experts
However, MoE models are quite complex and have certain limitations. One key issue is that they are very sensitive to how they are set up, which means a lot of time and effort is needed to find the right settings (called hyperparameters). A major problem with MoE models is something called “posterior collapse”. This happens when one part of the model becomes much more dominant than the others. It’s like a “rich gets richer” situation, where the other parts become less effective, or a case where all parts of the model end up being the same, making them less useful.
Also, the implementation of MoE models poses a considerable challenge, primarily due to the intricacies associated with dynamic routing and load-imbalanced computation Gale et al. (2023). Existing hardware and software for deep learning, such as TPUs and XLA compilers, often demand static knowledge of tensor shapes, making MoE implementation on TPU challenging.
While GPU implementation offers more flexibility, sparse computation compatibility becomes a hurdle. Striking the right balance between fixing the size of each expert to facilitate efficient computation and maintaining model quality creates a tradeoff between information preservation and hardware efficiency. This tradeoff, in turn, necessitates careful consideration during hyperparameter tuning, adding a layer of complexity to the implementation of MoE models, potentially offsetting their advantages. Given the formidable challenges in MoE model implementation, it becomes almost inevitable for researchers and practitioners to resort to specialized tools and frameworks, such as Tutel Hwang et al. (2023) or Megablocks Gale et al. (2023)
Departing from the horizontal expansion characteristic of MoE models, our new approach, DUS, introduces a vertical dimension. This method takes a different, less complex route compared to MoE. While MoE contends with horizontal complexities, DUS operates on a vertical plane, simplifying the process by just copying, trimming the front and back layers, and then reassembling them. This shift in approach offers a unique and more straightforward way of working, moving away from conventional MoE challenges.
Aside from its simple implementation, unlike MoE, which may face challenges such as posterior collapse, DUS offers a notable benefit in that there are fewer situations where its usage becomes devoid of meaning. This point highlights that DUS is a useful and dependable choice, demonstrating its strength and adaptability in different situations.

### 29 Dec 2023 v2

@@ Appendix A Contributions
  • Introduction of the SOLAR 10.7 Billion-Parameter Model: We have released the SOLAR 10.7B model, which is not only depthwise scaled but also continually pretrained. The availability of SOLAR 10.7B under the Apache 2.0 license permits commercial usage, enabling the integration of this advanced model into a diverse range of products and services. This bridges the gap between academic research and practical applications, fostering wider accessibility and utility in various fields.
@@ B.2 Mixture of Experts
However, efficient implementation of MoE models poses a considerable challenge, primarily due to the intricacies associated with dynamic routing and load-imbalanced computation Gale et al. (2023). Existing hardware and software for deep learning, such as TPUs and XLA compilers, often demand static knowledge of tensor shapes, making MoE implementation on TPU challenging.
While GPU implementation offers more flexibility, sparse computation compatibility becomes a hurdle. Striking the right balance between fixing the size of each expert to facilitate efficient computation and maintaining model quality creates a tradeoff between information preservation and hardware efficiency. This tradeoff, in turn, necessitates careful consideration during hyperparameter tuning, adding a layer of complexity to the implementation of MoE models, potentially offsetting their advantages. Given the formidable challenges in MoE model implementation, it becomes almost inevitable for researchers and practitioners to resort to specialized tools and frameworks, such as Tutel Hwang et al. (2023) or Megablocks Gale et al. (2023).
Departing from the horizontal expansion characteristic of MoE models, the DUS method introduces model scaling in the vertical dimension. Notably, DUS does not introduce dynamism in the scaled model, which significantly reduces the complexity when compared to MoE. This shift in approach offers a unique and more straightforward way of working, moving away from conventional MoE challenges. Not only that, DUS also undergoes continued pretraining to quickly recover performance of the scaled model.